# Central galaxy alignments

### v1

@@ 1 Introduction
Several observational works have investigated these so-called intrinsic alignments, which are expected to have important implications for the extraction of cosmological information from upcoming galaxy surveys (e.g. Hirata et al. 2007; Troxel & Ishak 2015; Joachimi et al. 2015; Kirk et al. 2015). Evidence suggests that the alignment of galaxies with each other and with large-scale cosmic structures is influenced not only by their luminosity, colour and star formation history, but also their position within the host halo and cosmic environment. In this context, red/elliptical galaxies in general tend to display more significant alignments on a variety of scales (see e.g. Sales & Lambas, 2004; Yang et al., 2006; Agustsson & Brainerd, 2010; Kirk et al., 2015; Rodriguez et al., 2022; Smith et al., 2023; Desai & Ryden, 2022). Moreover, red satellite galaxies show a stronger preference to align with the galactic plane of red central galaxies (Yang et al., 2005; Wang et al., 2008; Kiessling et al., 2015; Kirk et al., 2015; Libeskind et al., 2015; Welker et al., 2018; Pawlowski, 2018; Johnston et al., 2019). This specific connection between star formation activity, or lack thereof, and galaxy alignments is sometimes referred to as anisotropic quenching or angular conformity (Wang et al., 2008; Stott, 2022; Ando et al., 2023), an obvious reference to the correlation effect called galactic conformity (see Bray et al., 2016; Otter et al., 2020; Maier et al., 2022; Lacerna et al., 2022, and references therein).
The shape of spiral galaxies, on the other hand, is usually linked to their angular momentum, which has been proposed to emerge from the torque produced by the external gravitational field (Heavens et al., 2000; Catelan et al., 2001; Codis et al., 2015). Despite this connection, there is little observational evidence of alignments of the shapes of spiral galaxies (Zjupa et al., 2020; Johnston et al., 2019; Samuroff et al., 2023).
@@
The results of Rodriguez et al. (2022, 2023) prompted the additional analysis of Rodriguez et al. (2024), where the evolution of the intrinsic alignments was investigated for red and blue galaxies separately, in connection with their different assembly and merger histories. Mergers are in fact shown to be a major contributor to the buildup of these alignments for red central galaxies, as opposed to blue central galaxies, which are younger, experience fewer mergers, and have only recently acquired their oblate shape. The work of Rodriguez et al. (2024) emphasizes even more the multi-scale nature of the intrinsic alignments. Mergers tend to increase the alignment between the stellar content of central galaxies and their dark-matter halos, whereas the preferred orientation of halos with respect to the large-scale environment is initially stronger and tends to diminish with time.
@@ 2 Data
2.1 The IllustrisTNG hydrodynamical simulation
In this study, we utilize the galaxy and dark-matter halo catalogues from the IllustrisTNG300 simulation at ( $z=0$ ) (referred to as TNG300, Nelson et al. 2019; Pillepich et al. 2018). The IllustrisTNG magneto-hydrodynamical cosmological simulations are executed using the arepo moving-mesh code (Springel, 2010) and represent an enhanced iteration of the original Illustris simulations (Vogelsberger et al., 2014b, a; Genel et al., 2014). These simulations incorporate sub-grid models that address factors such as radiative metal-line gas cooling, star formation, chemical enrichment from SNII, SNIa, and AGB stars, as well as stellar feedback processes, formation of supermassive black holes with multi-mode quasar activity, and kinetic feedback from black holes. It adopts a cubic box with a side length of $205\,h^{-1}$ Mpc and is run with 25003 dark-matter particles, each with a mass of $4.0\times 10^{7}h^{-1}{\rm M_{\odot}}$. The initial conditions include 25003 gas cells, with each gas cell having a mass of $7.6\times 10^{6}h^{-1}{\rm M_{\odot}}$.
A friends-of-friends (FOF) algorithm is employed to identify dark matter haloes—referred to as groups in this paper—using a linking length of 0.2 times the mean inter-particle separation (Davis et al., 1985). The gravitationally bound substructures, hereafter referred to as subhaloes, are identified using the SUBFIND algorithm (Springel et al., 2001; Dolag et al., 2009). We define the stellar mass of galaxies, denoted as ${\rm M_{\ast}{}}$, as the total mass of all stellar particles associated with each subhalo. In our analysis, following the selection criteria used in many previous works that take simulation resolution into account, we consider all simulated galaxies with stellar mass greater than $10^{8.5}$ M, resulting on a total number of 429982 galaxies. In this study, we analyze central galaxies identified by a specific flag in the simulation, which have an absolute magnitude brighter than -19.5 in the r-band.
@@ 2.2 SDSS
Figure 1: Mass distribution of the halos in TNG300. The total sample is shown in a black solid line while the green, red, pink and light blue histograms represent the halos in filaments, clusters, cluster outskirts, and others, respectively.
@@
To identify galaxy groups in the SDSS DR18, we follow the procedure described by Rodriguez & Merchán (2020b). This algorithm combines the Friends-of-Friends (FoF) method (Merchán & Zandivarez, 2005) and halo-based techniques (Yang et al., 2005). The algorithm begins by detecting gravitationally bound systems through a percolation method adapted from Huchra & Geller (1982). Each group is assumed to contain at least one bright central galaxy, and its initial properties are estimated based on the total luminosity of its members. In the subsequent iterative step, the halo-based approach refines group memberships and recalculates halo properties on the basis of updated luminosity estimates. This process continues until convergence is reached, ensuring a reliable identification of systems across a broad range of group sizes, from small associations to rich clusters. The resulting group catalogue provides essential properties, including galaxy membership, spatial positions, and halo mass estimations ($M_{\text{group}}$), derived using an abundance matching technique (Vale & Ostriker, 2004; Kravtsov et al., 2004; Conroy et al., 2006; Behroozi et al., 2010). This method assumes a one-to-one relationship between a group’s characteristic luminosity and its dark-matter halo mass. Additionally, galaxies within each group are classified as central and satellites, with the brightest galaxy designated as the central galaxy, while the remaining members are considered satellites (Rodriguez et al., 2021). This catalogue demonstrates strong consistency with weak gravitational lensing mass estimates (Gonzalez et al., 2021) and effectively reproduces the observed central-satellite galaxy population properties seen in numerical simulations. Our final data set consists of galaxies with spectroscopic magnitudes, redshifts, angular positions, and group and halo properties, providing a robust foundation for analyzing the alignment of galaxies across different cosmic environments.
@@ 2.3 Large scale environment definitions
This work aims to study the alignment of central galaxies located in different components of the cosmic web. We identify the latter using the Discrete Persistent Structures Extractor (DisPerSE; Sousbie, 2011; Sousbie et al., 2011) structure finder. DisPerSE detects persistent topological features such as peaks, voids, walls, and particularly filamentary structures using discrete Morse theory (see Morse, 1934), operating under the assumption that a mathematical framework, known as the Morse complex, is an adequate description of the cosmic web. In this context, a set of manifolds can be related to the aforementioned cosmic web structures. In a nutshell, this algorithm identifies the critical points of the density field, which correspond to locations where the field’s gradient vanishes—namely, the maximum, minimum, or saddles of the density field. To determine the density field from the input particle distribution (galaxies in the present case), the algorithm employs the Delaunay tessellation field estimator (DTFE; see van de Weygaert & Schaap, 2009). For further details on the procedure and technical aspects, visit DisPerSE’s website.
@@
Figure 1 illustrates the mass distributions of halos in the different cosmic web environments defined above, specifically those of central galaxies with an r-band brightness greater than -19.5. These are the central galaxies that we will focus on throughout this work. By definition, the halos located in clusters occupy the most massive regions. In contrast, the other three environments show a similar mass range, with only minor differences in the average values of their distributions. In terms of the number of objects, the majority of halos are found in the environment categorized as others, followed by those in filaments, then on the cluster outskirts and finally in clusters. This distribution pattern fits the volumes occupied by each of these environments, which generally follow the same order. This distribution differs from that found in earlier studies (e.g., Ganeshaiah Veena et al., 2019), as we only consider bright central galaxies in our analysis.
@@
To define large-scale environments in SDSS DR18 in a manner analogous to those used in simulations and compare our results, we combine the group catalogue obtained using the method of Rodriguez & Merchán (2020a) with the DisPerSE implementation of Malavasi et al. (2020a). We selected central galaxies brighter than -19.5 in the r-band. This approach allowed us to maintain consistent definitions as before: clusters were identified using the same mass threshold based on the mass provided by the group identification method; the remaining environments were defined following the same guidelines. It is important to note that this environment identification was conducted in a redshift space, which could introduce additional uncertainties.
@@ 3.1 Simulation
In this work, we will study the alignments of the central galaxies, i.e. those associated with the brightest/massive galaxy of each group (BGG). We will take as reference the directions corresponding to the main axes of the inertia tensor ($I_{i,j}$) of the BGGs, defined as:
@@
where $i$ and $j$ correspond to the three spatial axes of the simulated box, i.e., $i,j=1,2,3$. The term $m_{n}$ represents the mass of the $n$-th particle, while $x_{i}^{n}$ and $x_{j}^{n}$ denote the positions of the $n$-th cell or particle along the $i$-th and $j$-th axes, respectively. These particle positions are measured relative to the center of the subhalo to which the particle belongs, defined as the position of the particle with the minimum gravitational potential energy. From these directions, we will define three angles that will provide insight into the internal alignment of the BGG and its alignment with the other group members, referred to as satellites.
We begin by introducing the first angle, which quantifies the alignment of the internal components of the BGG. In this approach, we use the angle between the principal shape axes corresponding to dark matter and stars, which we refer to as $\theta_{int}$, and is illustrated in the first diagram of Figure 2. Let ${\bf r}_{a}$, ${\bf r}_{b}$, and ${\bf r}_{c}$ be the three normalized eigenvectors of $I_{i,j}$ corresponding to the major, intermediate, and minor axes, respectively. Then $\theta_{int}$ is defined as:
@@ 4 Dependence of alignments on mass
In the left panel of Figure 3, we observe that the angle $\theta_{int}$ decreases as the mass increases for both the major and the minor axes. The minor axis exhibits values lower than the major axis for masses below $\sim 10^{14}M_{\odot}$, while both axes show similar values for masses above this threshold. This angle represents the alignment between the stars and the dark matter halo they inhabit, indicating that the major and minor axes of both the stars and the dark matter halo tend to become more aligned as the mass of the halo increases. Additionally, it suggests that for lower mass haloes, the minor axes are more closely aligned with each other than their respective major axes. The noticeable difference in alignment between the major and minor axes of galaxies residing in less massive halos may be attributed to their tendency to be spiral (oblate) (see, for example, Rodriguez et al., 2024). This morphology results in a more clearly defined minor axis compared to the other two axes, which are quite similar and lie within the same plane. In the most massive end, the predominant morphology is triaxial, meaning that the three axes are clearly defined, with no differences in alignment between the minor and major axes. This behavior of internal alignment may arise from the fact that the processes involved in the formation and evolution of galaxies vary depending on their mass. This variation underscores the complex nature of galaxy development, highlighting how mass plays a crucial role in shaping galaxies.
@@
The central panel of Figure 3 illustrates the relationship between $\theta_{sat}$ and mass. It is evident that this angle, measured with respect to the major (resp. minor) axis, shows a strong alignment (resp. misalignment) across the entire range of masses. Examining the dark matter estimates (black curves), we observe that unlike many other halo properties, $\theta_{sat}$ does not show a strong dependency on mass. Only for lower masses, around $\sim 10^{12}M_{\odot}$, alignment in both axes seems to be a small feature. It’s important to note that this angle is calculated for each satellite individually. As a result, at the lower mass end, there are many groups with only a few members. In contrast, at the higher mass end, there are fewer clusters but a greater number of satellites. Estimating the axes of the central galaxy using stellar material yields measurements closer to 60 degrees. Here, the alignment of the major axis with mass increased, but the minor axis showed the opposite trend. This behaviour is expected since, as illustrated in the left panel described above, the stars are misaligned in relation to their own dark matter halo, and this misalignment decreases as mass increases. As a result, measurements derived from dark matter and stars tend to converge and resemble each other as mass increases. The results of $\theta_{sat}$ indicate that satellite galaxies are preferentially distributed around the major axis.
@@
We finally focus on the angle between the central galaxy’s shape with that of the host group, $\theta_{ch}$, in the right panel of Fig. 3. Is it clear that this angle significantly decreases with mass, showing a progressive alignment between the shape of the galaxy and its associated group as mass increases. This trend is observed for both the major and minor axes when analyzing the shapes of dark matter and stellar components. Notably, measurements taken from stars exhibit greater misalignment than those related to dark matter. This discrepancy may result from the internal misalignment represented by $\theta_{int}$, indicating that the dark matter of the BGG is more closely aligned with the shape of the host halo than the stars of the BGG (see Figures 7 and 8 in Rodriguez et al., 2023).
@@
Previous studies have shown that central galaxies are aligned with their surroundings, and this alignment is more pronounced when considering dark matter instead of just stars. In this section, we analyzed how these alignments depend on mass and found that both the internal orientation angle $(\theta_{int}$) and the orientation angle of the central galaxy shape with respect to that of the host halo ($\theta_{ch}$) increase with mass. In contrast, the distribution of satellites ($\theta_{sat}$) does not exhibit a significant dependence on mass.

### v2

@@ 1 Introduction
Several observational works have investigated these so-called intrinsic alignments, which are expected to have important implications for the extraction of cosmological information from upcoming galaxy surveys (e.g. (Hirata et al., 2007; Troxel & Ishak, 2015; Joachimi et al., 2015; Kirk et al., 2015). Evidence suggests that the alignment of galaxies with each other and with large-scale cosmic structures is influenced not only by their luminosity, colour and star formation history, but also their position within the host halo and cosmic environment. In this context, red/elliptical galaxies in general tend to display more significant alignments on a variety of scales (see e.g. Sales & Lambas, 2004; Yang et al., 2006; Agustsson & Brainerd, 2010; Kirk et al., 2015; Rodriguez et al., 2022; Smith et al., 2023; Desai & Ryden, 2022). Moreover, red satellite galaxies show a stronger preference to align with the galactic plane of red central galaxies (Yang et al., 2005; Wang et al., 2008; Kiessling et al., 2015; Kirk et al., 2015; Libeskind et al., 2015; Welker et al., 2018; Pawlowski, 2018; Johnston et al., 2019). This specific connection between star formation activity, or lack thereof, and galaxy alignments is sometimes referred to as anisotropic quenching or angular conformity (Wang et al., 2008; Stott, 2022; Ando et al., 2023), an obvious reference to the correlation effect called galactic conformity (see Bray et al., 2016; Otter et al., 2020; Maier et al., 2022; Lacerna et al., 2022, and references therein).
The shape of spiral galaxies, on the other hand, is usually linked to their angular momentum, which has been proposed to arise from torques produced by the external gravitational field (Heavens et al., 2000; Catelan et al., 2001; Codis et al., 2015), and tends to align parallel to filaments, in contrast to the angular momentum of higher-mass galaxies, which is preferentially perpendicular (Welker et al. (2020); Barsanti et al. (2022); Kraljic et al. (2021), and e.g. Ganeshaiah Veena et al. (2019, 2021) for studies in simulations). Despite this connection, there is little observational evidence for shape alignments in spiral galaxies (Zjupa et al., 2020; Johnston et al., 2019; Samuroff et al., 2023). However, some studies based on observations and simulations indicate that central disk galaxies show greater misalignment with their host haloes, and that this misalignment decreases with increasing halo mass and ex situ stellar mass fraction (Xu et al., 2023a, b).
@@
The results of Rodriguez et al. (2022, 2023) prompted the additional analysis of Rodriguez et al. (2024), where the evolution of the intrinsic alignments was investigated for red and blue galaxies separately, in connection with their different assembly and merger histories. Mergers are in fact shown to be a major contributor to the buildup of these alignments for red central galaxies, as opposed to blue central galaxies, which are younger, experience fewer mergers, and have only recently acquired their oblate shape. This aligns with findings from other cosmological simulations, such as those in Lagos et al. (2018), based on the EAGLE simulations, which demonstrate the broader prevalence of merger-induced alignment effects. The work of Rodriguez et al. (2024) emphasizes even more the multi-scale nature of the intrinsic alignments. Mergers tend to increase the alignment between the stellar content of central galaxies and their dark-matter halos, whereas the preferred orientation of halos with respect to the large-scale environment is initially stronger and tends to diminish with time.
@@ 2 Data
2.1 The TNG hydrodynamical simulation
In this study, we utilize the galaxy and dark-matter halo catalogues from the Illustris TNG300 simulation at ( $z=0$ ) (referred to as TNG300, Nelson et al. 2019; Pillepich et al. 2018b). The TNG magneto-hydrodynamical cosmological simulations are executed using the arepo moving-mesh code (Springel, 2010) and represent an enhanced iteration of the original Illustris simulations (Vogelsberger et al., 2014b, a; Genel et al., 2014). These simulations incorporate sub-grid models that address factors such as radiative metal-line gas cooling, star formation, chemical enrichment from SNII, SNIa, and AGB stars, as well as stellar feedback processes, formation of supermassive black holes with multi-mode quasar activity, and kinetic feedback from black holes. It adopts a cubic box with a side length of $205\,h^{-1}$ Mpc and is run with 25003 dark-matter particles, each with a mass of $4.0\times 10^{7}h^{-1}{\rm M_{\odot}}$. The initial conditions include 25003 gas cells, with each gas cell having a mass of $7.6\times 10^{6}h^{-1}{\rm M_{\odot}}$.
A friends-of-friends (FOF) algorithm is employed to identify dark matter haloes—referred to as groups in this paper—using a linking length of 0.2 times the mean inter-particle separation (Davis et al., 1985). The gravitationally bound substructures, hereafter referred to as subhaloes, are identified using the SUBFIND algorithm (Springel et al., 2001; Dolag et al., 2009). We define the stellar mass of galaxies, denoted as ${\rm M_{\ast}{}}$, as the total mass of all stellar particles associated with each subhalo. In our analysis, following the selection criteria used in many previous works that take simulation resolution into account (e.g. Pillepich et al., 2018a; Rodriguez et al., 2023; Perez et al., 2024; Rodriguez et al., 2024), we consider all simulated galaxies with stellar mass ($M_{\star}$) greater than $10^{8.5}$ M, resulting on a total number of 429982 galaxies. In this study, we analyze central galaxies identified as those subhalos whose index matches the GroupFirstSub entry of their corresponding FoF group in the TNG300 simulation, selecting those with $M_{r}<-19.5$ in the $r$-band. This yields a final sample of 89,534 systems, each containing more than 880 dark matter particles and 100 stellar particles. The magnitude threshold matches the group-definition criterion from Rodriguez & Merchán (2020b), where groups must contain at least one galaxy brighter than $M_{r}=-19.5$ (defining the faintest observable centrals in optical surveys).
@@ 2.2 SDSS
Figure 1: Mass distributions of halos and their central galaxies in TNG300. The top panel shows the distribution of total halo masses ($M_{h}$), while the bottom panel displays the stellar mass ($M_{\star}$) distribution of central galaxies. In both panels, the total sample is represented by a solid black line, while different environmental subsets are indicated as follows: filaments are shown in green, clusters in red, cluster outskirts in pink, and others in light blue.
@@
To identify galaxy groups in the SDSS DR18, we follow the procedure described by Rodriguez & Merchán (2020b). This algorithm combines the Friends-of-Friends (FoF) method (Merchán & Zandivarez, 2005) and halo-based techniques (Yang et al., 2005). The algorithm begins by detecting gravitationally bound systems through a percolation method adapted from Huchra & Geller (1982). Each group is assumed to contain at least one bright central galaxy, and its initial properties are estimated based on the total luminosity of its members. In the subsequent iterative step, the halo-based approach refines group memberships and recalculates halo properties on the basis of updated luminosity estimates. This process continues until convergence is reached, ensuring a reliable identification of systems across a broad range of group sizes, from small associations to rich clusters. The resulting group catalogue provides essential properties, including galaxy membership, spatial positions, and halo mass estimations ($M_{\text{group}}$), derived using an abundance matching technique (Vale & Ostriker, 2004; Kravtsov et al., 2004; Conroy et al., 2006; Behroozi et al., 2010). This method assumes a one-to-one relationship between a group’s characteristic luminosity and its dark-matter halo mass. Additionally, galaxies within each group are classified as central and satellites, with the brightest galaxy designated as the central galaxy, while the remaining members are considered satellites (Rodriguez et al., 2021). This catalogue demonstrates strong consistency with weak gravitational lensing mass estimates (Gonzalez et al., 2021) and effectively reproduces the observed central-satellite galaxy population properties seen in numerical simulations, focusing primarily on the variations in halo occupancy distribution and the properties of member galaxies in different environments (e.g., Alfaro et al., 2022; Rodríguez-Medrano et al., 2023). Our final data set consists of galaxies with spectroscopic magnitudes, redshifts, angular positions, and group and halo properties, providing a robust foundation for analyzing the alignment of galaxies across different cosmic environments.
@@ 2.3 Large scale environment definitions
This work aims to study the alignment of central galaxies located in different components of the cosmic web. We identify the latter using the Discrete Persistent Structures Extractor (DisPerSE; Sousbie, 2011; Sousbie et al., 2011) structure finder. DisPerSE  [^2] http://www2.iap.fr/users/sousbie/web/html/indexd41d.html detects persistent topological features such as peaks, voids, walls, and particularly filamentary structures using discrete Morse theory (see Morse, 1934), operating under the assumption that a mathematical framework, known as the Morse complex, is an adequate description of the cosmic web. In this context, a set of manifolds can be related to the aforementioned cosmic web structures. In a nutshell, this algorithm identifies the critical points of the density field, which correspond to locations where the field’s gradient vanishes—namely, the maximum, minimum, or saddles of the density field. To determine the density field from the input particle distribution (galaxies in the present case), the algorithm employs the Delaunay tessellation field estimator (DTFE; see van de Weygaert & Schaap, 2009).
@@
The top panel of Figure 1 illustrates the mass distributions of halos (i.e $M_{h}$ is the mass of the central subhalos provided by the TNG300 simulation), in the different cosmic web environments defined above, specifically those of central galaxies with an r-band brightness greater than -19.5. These are the central galaxies that we will focus on throughout this work. By definition, the halos located in clusters occupy the most massive regions. In contrast, the other three environments show a similar mass range, with only minor differences in the average values of their distributions. In terms of the number of objects, the majority of halos are found in the environment categorized as others, followed by those in filaments, then on the cluster outskirts and finally in clusters. This distribution pattern fits the volumes occupied by each of these environments, which generally follow the same order. This distribution differs from that found in earlier studies (e.g., Ganeshaiah Veena et al., 2019), as we only consider bright central galaxies in our analysis. The bottom panel of Figure 1 shows the stellar mass distribution of the central galaxies in the halos, whose (halo) mass distributions were displayed in the top panel. It can be seen that the stellar mass distributions are similar to those shown for the halo mass across the different environments, albeit with greater scatter. This reflects a well-established fact: beyond the processes governing halo mass growth, internal astrophysical processes also shape the stellar mass of the central galaxy. This is the reason for the blurring of boundaries that are otherwise clear when considering halo mass alone. In other words, even when imposing a lower or upper halo mass cut to define environments, this does not translate linearly into a selection on the stellar mass of central galaxies. Nevertheless, this does not imply that the significant differences between environments are driven by sampling distinct stellar mass distributions. Therefore, throughout the analysis in this work we will focus on halo mass.
@@
To define large-scale environments in SDSS DR18 in a manner analogous to those used in simulations and compare our results, we combine the group catalogue obtained using the method of Rodriguez & Merchán (2020a) with the DisPerSE implementation of Malavasi et al. (2020a). We selected central galaxies brighter than -19.5 in the r-band. This approach allowed us to maintain consistent definitions as before: clusters were identified using the same mass threshold based on the mass provided by the group identification method; the remaining environments were defined following the same guidelines. It is important to note that this environment identification in redshift space introduces systematic effects like the Fingers-of-God elongation and line-of-sight projection artefacts. These effects may dominate the uncertainty budget in environment classification, particularly for cluster regions and filament connectivity.
@@ 3.1 Simulation
In this work, we investigate the alignment properties of central galaxies, focusing on the brightest/most massive group galaxies (BGGs). Building upon our previous methodological developments (Rodriguez et al., 2023, 2024) and following established best practices for shape measurement (Zemp et al., 2011; Bassett & Foster, 2019), we analyze these systems using exclusively stellar and dark matter particles enclosed within twice the half-dark matter mass radius. We compute the inertia tensor ($I_{i,j}$) of the BGGs as:
@@
where $i$ and $j$ correspond to the three spatial axes of the simulated box, i.e., $i,j=1,2,3$. The term $m_{n}$ represents the mass of the $n$-th, while $x_{i}^{n}$ and $x_{j}^{n}$ denote the positions of the $n$-th particle along the $i$-th and $j$-th axes, respectively. These particle positions are measured relative to the centre of the subhalo to which the particle belongs, defined as the position of the particle with the minimum gravitational potential energy. Let $\mathbf{r}_{a}$, $\mathbf{r}_{b}$, and $\mathbf{r}_{c}$ be the three normalized eigenvectors of $I_{i,j}$ corresponding to the major, intermediate, and minor axes, respectively.
From these directions, we will define three angles that will provide insight into the internal alignment of the BGG and its alignment with the other group members, referred to as satellites.
We begin by introducing the first angle, which quantifies the alignment of the internal components of the BGG. In this approach, we use the angle between the principal shape axes corresponding to dark matter and stars, which we refer to as $\theta_{int}$, is illustrated in the first diagram of Figure 2 and is defined as:
@@ 4 Dependence of alignments on mass
In the left panel of Figure 3, we observe that the angle $\theta_{int}$ decreases as the mass increases for both the major and the minor axes. The minor axis exhibits values lower than the major axis for masses below $\sim 10^{14}M_{\odot}$, while both axes show similar values for masses above this threshold. This angle represents the alignment between the stars and the dark matter halo they inhabit, indicating that the major and minor axes of both the stars and the dark matter halo tend to become more aligned as the mass of the halo increases. Additionally, it suggests that for lower mass haloes, the minor axes are more closely aligned with each other than their respective major axes. The noticeable difference in alignment between the major and minor axes of galaxies residing in less massive halos may be attributed to their tendency to be spiral (oblate) (see, for example, Rodriguez et al., 2024). This morphology results in a more clearly defined minor axis compared to the other two axes, which are quite similar and lie within the same plane. In the most massive end, the predominant morphology is triaxial, meaning that the three axes are clearly defined, with no differences in alignment between the minor and major axes. This behavior of internal alignment may arise from the fact that the processes involved in the formation and evolution of galaxies vary depending on their mass. This variation underscores the complex nature of galaxy development, highlighting how halo mass plays a crucial role in shaping galaxies.
@@
The central panel of Figure 3 illustrates the relationship between $\theta_{sat}$ and mass. It is evident that this angle, measured with respect to the major (resp. minor) axis, shows a strong alignment (resp. misalignment) across the entire range of masses. Examining the dark matter estimates (black curves), we observe that unlike many other halo properties, $\theta_{sat}$ does not show a strong dependency on mass. Only for lower masses, around $\sim 10^{12}M_{\odot}$, alignment in both axes seems to be a small feature. It’s important to note that this angle is calculated for each satellite individually. As a result, at the lower halo mass end, there are many groups with only a few members. In contrast, at the higher mass end, there are fewer clusters but a greater number of satellites. Estimating the axes of the central galaxy using stellar material yields measurements closer to 60 degrees. Here, the alignment of the major axis with mass increased, but the minor axis showed the opposite trend. This behaviour is expected since, as illustrated in the left panel described above, the stars are misaligned in relation to their own dark matter halo, and this misalignment decreases as halo mass increases. As a result, measurements derived from dark matter and stars tend to converge and resemble each other as halo mass increases. The results of $\theta_{sat}$ indicate that satellite galaxies are preferentially distributed around the major axis.
@@
We finally focus on the angle between the central galaxy’s shape with that of the host group, $\theta_{ch}$, in the right panel of Fig. 3. Is it clear that this angle significantly decreases with mass, showing a progressive alignment between the shape of the galaxy and its associated group as halo mass increases. This trend is observed for both the major and minor axes when analyzing the shapes of dark matter and stellar components. Notably, measurements taken from stars exhibit greater misalignment than those related to dark matter. This discrepancy may result from the internal misalignment represented by $\theta_{int}$, indicating that the dark matter of the BGG is more closely aligned with the shape of the host halo than the stars of the BGG (see figures 7 and 8 in Rodriguez et al., 2023).
@@
Previous studies have shown that central galaxies are aligned with their surroundings, and this alignment is more pronounced when considering dark matter instead of just stars. In this section, we analyzed how these alignments depend on halo mass and found that both the internal orientation angle $(\theta_{int}$) and the orientation angle of the central galaxy shape with respect to that of the host halo ($\theta_{ch}$) increase with mass. In contrast, the distribution of satellites ($\theta_{sat}$) does not exhibit a significant dependence on mass.